OpenStudy (anonymous):

Find 87^2 using mental math

OpenStudy (anonymous):

still need help on what

OpenStudy (anonymous):

87*87

OpenStudy (anonymous):

yeah

OpenStudy (p0sitr0n):

87 x 87 = (90 -3) x (90 -3) = 90 ^2 -2 x 90 x 30 + 9 = 8100 - 540 + 9 = 8100 - 531 = 7569

OpenStudy (anonymous):

=7569

OpenStudy (kaptain_mittens):

p0sitr0n can you help me now por favor

OpenStudy (p0sitr0n):

yep yep

OpenStudy (anonymous):

thankyou so much everyone as you can tell I stink at algebra

OpenStudy (anonymous):

mental math can be hard